# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 7 Results
For the LLM-based social network model, we conducted 10 distinct runs, constrained by OpenAI API limitations; the current simulation, with $N=1000$ and $T=30$, required approximately 100,000 GPT-3.5 API calls to generate experimental and ground truth results. Each run employs a unique treatment allocation following a staggered rollout design across three stages with $\vec{p}=(0.2,0.5,0.8)$, each spanning 10 periods. This design implies that on average 20% of units received the intervention in the initial 10 periods, followed by an additional 30% in the subsequent 10 periods, and so forth.
@@
For the remaining five experiments, we conducted 100 independent runs for each setting, utilizing a fresh treatment allocation for each run through a staggered rollout design. The design comprises four equal-length stages with treatment probabilities $\vec{p}=(0,0.2,0.4,0.6)$. In each figure’s leftmost panel, we display the temporal evolution of outcomes through their mean and standard deviation, along with the 95th percentile across runs.
@@
Figure 15: Ascending auction model with $N=500$ objects.
The second panels of Figures 11-18 display the box plot of the average total treatment effect (TTE) across multiple time periods. The TTE contrasts the counterfactual of all units under treatment against all units under control:
In each setting, we carefully select $L$ so that the TTE in (14) covers all periods with nonzero treatment probability, ensuring our benchmark estimators remain meaningful. The results compare ground truth (GT) values against estimates obtained from three methods: our proposed causal message-passing approach (CMP), the difference-in-means estimator (DM), and the Horvitz-Thompson estimator (HT) [^8] Difference-in-means (DM) and Horvitz-Thompson (HT) are expressed as: $\hat{\tau}_{\text{DM}}:=\frac{1}{L}\sum_{t=T-L+1}^{T}\Big{(}\frac{\sum_{i=1}^{N}Y^{i}_{t}{}W^{i}_{t}}{\sum_{i=1}^{N}W^{i}_{t}}-\frac{\sum_{i=1}^{N}Y^{i}_{t}{}(1-W^{i}_{t})}{\sum_{i=1}^{N}(1-W^{i}_{t})}\Big{)},\quad\quad\hat{\tau}_{\text{HT}}:=\frac{1}{LN}\sum_{t=T-L+1}^{T}\sum_{i=1}^{N}\left(\frac{Y^{i}_{t}{}W^{i}_{t}}{\mathbb{E}[W^{i}_{t}]}-\frac{Y^{i}_{t}{}(1-W^{i}_{t})}{\mathbb{E}[1-W^{i}_{t}]}\right).$ (Sävje et al., 2021). Finally, the rightmost two panels in each figure display the CFE under the ground truth and CMP estimates for all-control and all-treatment conditions, along with their respective 95th percentiles.
In implementing Algorithm 2, we employ five validation batches ($b_{v}=5$). To select candidate estimators, we begin with a base model where each outcome is expressed as a linear function of two components: the sample mean of outcomes from the previous round and the current treatment allocation means. We then systematically modify this model by incorporating additional first-order and higher-order terms. The configurations also span batch counts from 200 to 2000 and batch sizes ranging from 0.1 to 20 percent of the population size. To estimate parameters, we employ Ridge regression with penalty parameters logarithmically spaced from $10^{-4}$ to $10^{4}$. These parameters are comprehensively combined to generate a diverse set of potential estimators, with time blocks aligned to experimental stages. For example, when $T=40$ and the design $\vec{p}=(0,0.2,0.4,0.6)$ with equal length blocks is used, $\mathfrak{L}_{\mathsf{tb}}$ has four elements, one corresponding to each block with a fixed treatment probability. Then, the selection process incorporates both domain knowledge and observed data characteristics. For instance, the pronounced temporal patterns evident in the left panels of Figures 16-18, observed in the New York City taxi model, exercise encouragement program, and data center model, necessitate estimators with detrending steps (see Remark 5.2). Computational efficiency is maintained by constraining the estimator search space based on the experimental context.
Overall, our framework demonstrates robust performance across all six scenarios, successfully estimating counterfactual evolutions despite strong seasonality patterns and without requiring information about the underlying interference network. The proposed method achieves significantly better performance than both DM and HT estimators, even in settings with subtle treatment effects. As illustrated in Figures 11-16, CMP yields estimates with both smaller bias and variance in different scenarios. The effectiveness of our method is particularly evident in the challenging scenarios presented in Figures 12-15 and 18, where conventional estimators struggle to reliably determine even the direction of treatment effects. These comprehensive results establish our framework’s capability to deliver precise estimates of counterfactual evolutions and treatment effects across diverse experimental settings.
@@
Figure 18: Data Center model with $N=1,000$ servers.
@@
Selecting a predetermined number of batches for a given batch size $n^{\mathcal{S}}$ presents a significant computational challenge, particularly in large-scale problems with time-varying treatment allocations across units. For staggered rollout designs, we implement a heuristic approach while deferring comprehensive analysis to future research. Our heuristic consists of three steps. First, we order units by their treatment duration, defined as the number of time periods under treatment. Second, we select two blocks of size $n^{\mathcal{S}}$—one that slides through the ordered list to cover all treatment durations, and another chosen randomly to ensure sufficient between-batch variation. Third, we select individual units from these merged blocks with equal probability to generate batches with average size $n^{\mathcal{S}}$. This procedure maintains computational efficiency while ensuring batches with diverse treatment allocations with high probability.
@@ 8 Concluding Remarks
This work presents several contributions to the study of causal inference under network interference: a distribution-preserving network bootstrap approach that enables resampling from networked populations with unobserved interaction patterns, a counterfactual cross-validation framework to validate estimation methods, and a benchmark toolbox containing six experimental environments. Our empirical results suggest the framework can be effective across diverse settings, showing promising performance compared to existing approaches when handling complex temporal patterns and unobserved network structures. Future work exploring advanced machine learning architectures and incorporating domain-specific knowledge may further enhance the framework’s practical utility while maintaining its theoretical properties.
Appendices Organization
@@
These appendices provide detailed statements and proofs of theoretical results that support the main body of the paper.
@@
The first appendix presents the technical results, beginning with essential notation and a reformulation of the outcome specification. We then establish the outcome decomposition rules through a sequence of proofs. This begins with Lemmas 9.1-9.3, culminating in the proof of Theorem 4.4 and Corollary 4.7. The discussion then progresses to a rigorous statement of batch-level state evolution in §9.4, which was informally introduced in Theorem 4.2. This is followed by a brief overview of the conditioning technique in §9.5, which is essential for proving the state evolution equation, as detailed in §9.6.
@@
Appendix 10 demonstrates how consistent estimation of state evolution parameters enables consistent estimation of desired counterfactuals. We present the necessary assumptions, provide and prove the main theoretical results, and detail the corresponding algorithms. We then analyze the application of these results to Bernoulli randomized designs, concluding in presenting two families of estimators in §10.2.1 and §10.2.2.
@@
Building on these results, §11 presents an extension to the causal message-passing methodology that addresses strong time-trends, enabling general counterfactual estimation even in the presence of seasonality or temporal patterns. The appendices conclude with §12, which presents auxiliary theorems necessary for our main proofs.
@@ 9.1 Notations
For any set $S$, the indicator function $\mathds{1}_{S}(\omega)$ evaluates to $1$ if $\omega$ belongs to $S$, and 00 otherwise. We define $\mathbb{R}^{n\times m}$ as the set of matrices with $n$ rows and $m$ columns. Given a matrix $\bm{M}$, we denote its transpose as $\bm{M}^{\top}$. Additionally, we represent a matrix of ones with dimensions $n\times m$ as $\bm{1}_{n\times m}\in\mathbb{R}^{n\times m}$. The symbol =$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}$ is used to denote equality in distribution, while =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;$ is used for equalities that hold almost surely with respect to the reference probability measure $\mathbb{P}$.
@@ 9.2 Preliminaries
Compared to (3), the specification in (15) and (16) is more comprehensive, incorporating both the complete outcome history and the full treatment allocation matrix. This generalization enables us to capture more complex temporal dynamics, including additional lag terms and potential anticipation effects of treatments. The functions $g_{t}$ and $h_{t}{}$ can also be specified to include only a finite number $l\in\mathbb{N}$ of historical lag terms. While all subsequent results hold for the general model in (15) and (16), for notational simplicity, we present the proofs using only one lag term $\vec{Y}_{t}{}(\bm{W}{})$. Furthermore, when the context is clear, we omit the treatment matrix notation $\bm{W}{}$ and write simply $\vec{Y}_{t}{}$ as the vector of outcomes at time $t$.
@@ 9.3 Outcome Decomposition Rule
Here, $[\vec{\text{e}}_{1}|\ldots|\vec{\text{e}}_{s}]^{\perp}$ represents the orthogonal complement of the first $s$ standard basis vectors. Then, a similar argument to the one in Step 1 implies that the matrix $\widetilde{\bf A}_{s}\in\mathbb{R}^{N\times(N-s)}$ has i.i.d. entries with a distribution of ${\cal N}(0,{\sigma^{2}}/{N})$. Furthermore, it yields that $\widetilde{\bf A}_{s}$ (and consequently ${\bf A}_{s}$, see Eq. (23)) is independent of the vector ${\bf A}_{s-1}\vec{V}_{s-1}{}={\bf A}({\bf I}_{N}-\bm{V}_{s-2}{}\bm{V}^{\top}_{s-2}{})\vec{V}_{s-1}{}={\bf A}\vec{V}_{s-1}{}$; this holds true because of Eq. (23) and the fact that $\vec{V}^{\top}_{j}{}\vec{V}_{s-1}{}=0$, for $j=0,\ldots,s-2$.
@@
By combining the results of Lemmas 9.1 and 9.2, we arrive at the conclusion of Lemma 9.3.
@@ 9.4 Batch-level State Evolution
Next, we analyze the large-sample behavior of the outcomes for a subpopulation of units. Specifically, let $\mathcal{S}\subset[N]$ represent an arbitrary subpopulation, with its size $\left|{\mathcal{S}}\right|$ increasing indefinitely as the population size $N$ grows large to infinity. We investigate the asymptotic behavior of the elements in $\bm{Y}{}$ as $N$ approaches infinity, This provides valuable insights into the evolution of outcomes within the subpopulation. {assumption} Fixing $T\in\mathbb{N}$ and $k\geq 2$, we assume that
@@
    For all $t\in[T]$, the function $g_{t}:\mathbb{R}^{1+T+M}\rightarrow\mathbb{R}$ is a ${{\cal C}{\cal P}}(\frac{k}{2})$ function.
@@
    For all $t\in[T]$, the function $h_{t}{}:\mathbb{R}^{1+T+M}\rightarrow\mathbb{R}$ is a ${{\cal C}{\cal P}}(1)$ function.
@@
Assumption 9.4 comprises a collection of regularity conditions on the model attributes. The first two parts ensure that the functions $g_{t}$ and $h_{t}{}$ do not demonstrate fast explosive behavior, thereby ensuring the well-posedness of the large system asymptotic. The final part pertains to the initial observation of the network and corroborates that the function $g_{0}$ is non-degenerate. This ensures that initial observations provide meaningful information and contribute to the evolution of outcomes. {assumption} Fix $T\in\mathbb{N}$, $k\geq 2$, and a subpopulation $\mathcal{S}\subset[N]$, where the size $\left|{\mathcal{S}}\right|$ grows to infinity as $N\rightarrow\infty$. We assume the following statements hold:
@@
    $p_{y}^{\mathcal{S}}(N)$ denotes the empirical distribution of the initial outcomes $Y^{i}_{0}{(N)}$ with $i\in\mathcal{S}$, and converges weakly to a probability measure $p_{y}^{\mathcal{S}}$ such that $\mathbb{E}_{p_{y}^{\mathcal{S}}}\big{[}{\|Y_{0}{}\|}^{k}\big{]}<\infty$,
@@
    $p_{x}^{\mathcal{S}}(N)$ denotes the empirical distribution of the covariate vectors ${\vec{X}}^{i}(N)$ with $i\in\mathcal{S}$, and converges weakly to a probability measure $p_{x}^{\mathcal{S}}$ such that $\mathbb{E}_{p_{x}^{\mathcal{S}}}\big{[}{\|{\vec{X}}\|}^{k}\big{]}<\infty$,
@@
    $p_{w}^{\mathcal{S}}(N)$ denotes the empirical distribution of the treatment vectors $\vec{W}^{i}_{N}$ with $i\in\mathcal{S}$, and converges weakly to a probability measure $p_{w}^{\mathcal{S}}$ such that $\mathbb{E}_{p_{w}^{\mathcal{S}}}\big{[}{\|\vec{W}\|}^{k}\big{]}<\infty$.
@@
    For all $i$ and any $t\in[T]_{0}$, let $p_{\mu^{i}}(N)$ and $p_{\mu_{t}^{i}}(N)$ be, respectively, the empirical distribution of the elements of the vectors $\vec{\mu}^{\;i\cdot}:=(\mu^{i1},\ldots,\mu^{iN})^{\top}$ and $\vec{\mu}^{\;i\cdot}_{t}:=(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})^{\top}$. Then, $p_{\mu^{i}}(N)$ converges weakly to $p_{\mu^{i}}$ and $p_{\mu_{t}^{i}}(N)$ converges weakly to $p_{\mu_{t}^{i}}$. Also, $\mathbb{E}_{p_{\mu^{i}}}\big{[}{\|{M}^{i}\|}^{k}\big{]}<\infty$ and $\mathbb{E}_{p_{\mu^{i}_{t}}}\big{[}{\|{M}_{t}^{i}\|}^{k}\big{]}<\infty$.
    In addition, the limit distributions ($p_{\mu^{i}}$ and $p_{\mu^{i}_{t}}$) are independent of other randomnesses in the model and if $\bar{\mu}^{i}$ denotes the mean of a random variable under probability measure $p_{\mu^{i}}$ (i.e., $\bar{\mu}^{i}=\mathbb{E}_{p_{\mu^{i}}}[{M}^{i}]$), the empirical distributions of $\bar{\mu}^{i},\;i\in\mathcal{S}$, denoted by $p_{\mu}^{\mathcal{S}}(N)$, converges weakly to a probability measure $p_{\mu}^{\mathcal{S}}$. Likewise, we let $p_{\mu_{t}}^{\mathcal{S}}$ denote the weak limit of the empirical distribution of the means under probability measures $p_{\mu^{i}_{t}}$. Finally, $\mathbb{E}_{p_{\mu}^{\mathcal{S}}}\big{[}{\|{M}^{\mathcal{S}}\|}^{k}\big{]}<\infty$ as well as $\mathbb{E}_{p_{\mu_{t}}^{\mathcal{S}}}\big{[}{\|{M}_{t}^{\mathcal{S}}\|}^{k}\big{]}<\infty$.

### v2

@@ 7 Results and Discussions
For the LLM-based social network, we conducted 10 distinct runs, constrained by OpenAI API limitations. Each run with $N=1,000$ and $T=31$ required approximately 100,000 GPT-3.5 API calls to generate experimental and ground truth results. Each run employs a unique treatment allocation following a staggered rollout design across three stages with $\vec{p}=(0.2,0.5,0.8)$, each spanning 10 periods. This design ensures that on average 20% of units receive intervention in the initial 10 periods, with an additional 30% in the subsequent 10 periods, and so forth. All runs begin with a single historical observation with no treatment to initialize the recursive estimation algorithm.
@@
For the remaining five experiments, we conducted 100 independent runs for each setting with treatment probabilities $\vec{p}=(0.1,0.2,0.5)$ across three stages (see Table 2 for complete details). In the leftmost panel in Figures 1925, we display the temporal trajectories of observed outcomes through their mean and standard deviation, along with the 95th percentile across runs.
The second panels of Figures 1925 display box plots of the average total treatment effect (TTE):
In each setting, we select $L$ so that the TTE in (19) covers all periods in the last experimental stage. The results compare ground truth (GT) values against estimates from four methods: our proposed causal message-passing approach (CMP), which employs DPNB and cross-validation; basic causal message-passing (bCMP), which extends the estimator of Shirani and Bayati, (2024) without DPNB or cross-validation; the difference-in-means estimator (DM); and the Horvitz-Thompson estimator (HT) [^7] Difference-in-means (DM) and Horvitz-Thompson (HT) are expressed as: $\hat{\tau}_{\text{DM}}:=\frac{1}{L}\sum_{t=T-L+1}^{T}\Big(\frac{\sum_{i=1}^{N}Y^{i}_{t}{}W^{i}_{t}}{\sum_{i=1}^{N}W^{i}_{t}}-\frac{\sum_{i=1}^{N}Y^{i}_{t}{}(1-W^{i}_{t})}{\sum_{i=1}^{N}(1-W^{i}_{t})}\Big),\quad\quad\hat{\tau}_{\text{HT}}:=\frac{1}{LN}\sum_{t=T-L+1}^{T}\sum_{i=1}^{N}\left(\frac{Y^{i}_{t}{}W^{i}_{t}}{\mathbb{E}[W^{i}_{t}]}-\frac{Y^{i}_{t}{}(1-W^{i}_{t})}{\mathbb{E}[1-W^{i}_{t}]}\right).$ (Sävje et al., 2021). Finally, the rightmost panels display the counterfactual evolution (CFE) under ground truth and CMP estimates for all-control and all-treatment conditions, along with their respective 95th percentiles across different runs.
@@
In implementing Algorithm 2, we select candidate estimators, starting with a base model where each outcome is expressed as a linear function of two components: the sample mean of previous round outcomes and the current treatment allocation means. We then systematically modify this model by incorporating higher-order and interaction terms. Then, the parameters in Table 2 are comprehensively combined to generate a diverse set of configurations, with time blocks aligned to experimental stages. For example, when $T=7$, $\mathfrak{L}_{\mathsf{tb}}$ has three elements: $([0,3],[3,5],[5,7])$.
Our framework demonstrates robust performance across all scenarios, successfully estimating counterfactual evolutions in the presence of strong seasonality patterns and without requiring information about the underlying interference network. As illustrated in Figures 1925, CMP yields estimates with smaller bias, closely tracking the ground truth in most scenarios. The effectiveness of our method is particularly evident in the challenging scenarios presented in Figures 2022, where conventional estimators struggle to reliably determine the direction of treatment effects. Meanwhile, our method successfully estimates the treatment effect despite its small signal magnitude.
These results demonstrate our framework’s ability to accurately estimate counterfactual evolutions and treatment effects across different experimental contexts. The robust performance, spanning social networks to distributed service systems like data centers, can empower decision makers with reliable causal insights when network interference undermines conventional approaches.
@@
Figure 25: Data Center model with $N=2,000$ servers.
Figure 26: Auction model with $N=200$ objects.
@@
Selecting a predetermined number of batches for a given batch size $\mathsf{n_{b}}$ presents a significant computational challenge, particularly in large-scale problems with time-varying treatment allocations across units. For staggered rollout designs, we implement a heuristic approach while deferring comprehensive analysis to future research. Our heuristic consists of three steps. First, we order units by their treatment duration, defined as the number of time periods under treatment. Second, we select two blocks of size $\mathsf{n_{b}}$: one that slides through the ordered list to cover all treatment durations, and another chosen randomly to ensure sufficient between-batch variation. Third, we select individual units from these merged blocks with equal probability to generate batches with average size $\mathsf{n_{b}}$. This procedure maintains computational efficiency while ensuring batches with diverse treatment allocations with high probability. For more details, see Appendix 9.
@@ 8 Concluding Remarks
This work presents several contributions to the study of causal inference under network interference: a distribution-preserving network bootstrap approach that enables resampling from networked populations with unobserved interaction patterns, a counterfactual cross-validation framework to validate estimation methods, and an interference gym containing six experimental environments. Our empirical results suggest the framework can be effective across diverse settings, showing promising performance when handling temporal patterns and complex unobserved network structures.
Future work could explore advanced machine learning architectures and incorporate domain-specific knowledge to further enhance the framework’s practical utility while maintaining its theoretical properties. An important direction involves handling scenarios with more complex time trends, particularly where temporal patterns are significantly altered by treatment allocation, which would extend the framework’s applicability to more challenging experimental contexts.
Appendices
@@
The following appendices provide supporting material for the main paper, including detailed theoretical statements, rigorous proofs, and technical derivations that establish the foundation for our results.
Organization
@@
The first appendix outlines our heuristic batching technique for the DPNB implementation. The second appendix presents the technical results, beginning with essential notation and a reformulation of the outcome specification. We then establish the outcome decomposition rules through a sequence of proofs. This begins with Lemmas 10.1-10.3, culminating in the proof of Theorem 4.4 and Corollary 4.5. The discussion then progresses to a rigorous statement of batch-level state evolution in §10.4, that a simpler version of which was informally introduced in Theorem 4.2. This is followed by a brief overview of the conditioning technique in §10.5, which is essential for proving the state evolution equation, as detailed in §10.6.
@@
Appendix 11 demonstrates how consistent estimation of state evolution parameters enables consistent estimation of desired counterfactuals. We present the necessary assumptions, prove the main theoretical results, and detail the corresponding algorithms. We then analyze the application of these results to Bernoulli randomized designs, concluding in presenting three families of estimators in §11.2.1, §11.2.2, and §11.3.
@@
Building on these results, §12 presents an extension to the causal message-passing methodology that addresses strong time-trends, enabling counterfactual estimation in the presence of seasonality or temporal patterns. The appendices conclude with §13, which presents auxiliary theorems necessary for our main proofs.
9 Heuristic Batching Technique for DPNB
To implement DPNB effectively, we must construct batches that exhibit sufficient variation in treatment exposure levels. The following pseudocode demonstrates our approach for generating such batches under staggered rollout designs, where treatment allocation probabilities evolve systematically over time.
Algorithm 3 Create Random Batches for DPNB Implementation
Batch size $\mathsf{s_{b}}$ and number of batches $\mathsf{n_{b}}$
Sort units by their treatment duration (number of time periods under treatment)
Create $\mathsf{n_{b}}$ evenly spaced positions along the sorted units
for each batch $i$ do
  Select a systematic block of size $\mathsf{s_{b}}$ starting at position $i$ in evenly spaced positions
  Select a random block of size $\mathsf{s_{b}}$ from anywhere in the dataset
  Merge the two blocks and remove duplicates to form candidate pool
  Calculate selection probability to achieve desired batch size on average
  Randomly select units from the candidate pool
end for
$\mathsf{n_{b}}$ distinct batches with average size of $\mathsf{s_{b}}$
Algorithm 3 works in three key steps. First, it orders all participants based on how long they’ve been receiving treatment. Second, it selects two groups of participants: one group that systematically slides through the ordered list to ensure coverage of all treatment durations, and another group chosen completely randomly to add variation. Finally, it combines these two groups and randomly selects individual participants from this pool with equal probability, creating batches that naturally contain diverse treatment allocations while maintaining computational efficiency, see Figure 27.
Step 1: Sort units by treatment duration11234455667788991010Step 2a: Select systematic block (slides through ordered list)11234455667788991010Step 2b: Select random block11234455667788991010Step 3a: Merge blocks to form candidate pool11234455667788991010Step 3b: Randomly select units from candidate pool11234455667788991010
Figure 27: Visualization of the Random Batch Creation Algorithm for Staggered Rollout Designs.
@@ 10.1 Notations
For any set $S$, the indicator function $\mathds{1}_{S}(\omega)$ evaluates to $1$ if $\omega$ belongs to $S$, and 0 otherwise. We define $\mathbb{R}^{n\times m}$ as the set of matrices with $n$ rows and $m$ columns. Given a matrix $\bm{M}$, we denote its transpose as $\bm{M}^{\top}$. Additionally, we represent a matrix of ones with dimensions $n\times m$ as $\bm{1}_{n\times m}\in\mathbb{R}^{n\times m}$. The symbol $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}$ is used to denote equality in distribution, while $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;$ is used for equalities that hold almost surely with respect to the reference probability measure $\mathbb{P}$.
@@ 10.2 Preliminaries
Compared to Eq. (3), the specification in Eqs. (20) and (21) is more comprehensive, incorporating both the complete outcome history and the full treatment allocation matrix. This generalization enables us to capture more complex temporal dynamics, including additional lag terms and potential anticipation effects of treatments. The functions $g_{t}$ and $h_{t}$ can also be specified to include only a finite number $l\in\mathbb{N}$ of historical lag terms. While all subsequent results hold for the general model in (20) and (21), for notational simplicity, we present the proofs using only one lag term $\vec{Y}_{t}{}(\bm{W}{})$. Furthermore, when the context is clear, we omit the treatment matrix notation $\bm{W}{}$ and write simply $\vec{Y}_{t}{}$ as the vector of outcomes at time $t$.
@@ 10.3 Outcome Decomposition Rule
Here, $\vec{\text{e}}_{1}|\ldots|\vec{\text{e}}_{s}]^{\perp}$ represents the orthogonal complement of the first $s$ standard basis vectors. Then, a similar argument to the one in Step 1 implies that the matrix $\widetilde{\bf A}_{s}\in\mathbb{R}^{N\times(N-s)}$ has i.i.d. entries with a distribution of ${\cal N}(0,{\sigma^{2}}/{N})$. Furthermore, it yields that $\widetilde{\bf A}_{s}$ (and consequently ${\bf A}_{s}$, see Eq. (28)) is independent of the vector ${\bf A}_{s-1}\vec{V}_{s-1}{}={\bf A}({\bf I}_{N}-\bm{V}_{s-2}{}\bm{V}^{\top}_{s-2}{})\vec{V}_{s-1}{}={\bf A}\vec{V}_{s-1}{}$; this holds true because of Eq. (28) and the fact that $\vec{V}^{\top}_{j}{}\vec{V}_{s-1}{}=0$, for $j=0,\ldots,s-2$.
@@
By combining the results of Lemmas 10.1 and 10.2, we arrive at the conclusion of Lemma 10.3.
@@ 10.4 Batch-level State Evolution
Next, we analyze the large-sample behavior of the outcomes for a subpopulation of units. Specifically, let $\mathcal{S}\subset[N]$ represent an arbitrary subpopulation [^8] This generalizes the result in §4 which focuses on treatment-allocation-based subpopulations., with its size $\left|{\mathcal{S}}\right|$ increasing indefinitely as the population size $N$ grows large to infinity. We investigate the asymptotic behavior of the elements in $\bm{Y}{}$ as $N$ approaches infinity. This provides valuable insights into the evolution of outcomes within the subpopulation. {assumption} Fixing $T\in\mathbb{N}$ and $k\geq 2$, we assume that
@@
    For all $t\in[T]$, the function $g_{t}:\mathbb{R}\times\mathbb{R}^{T+1}\times\mathbb{R}^{n^{\text{x}}}\rightarrow\mathbb{R}$ is a ${{\cal C}{\cal P}}(\frac{k}{2})$ function.
@@
    For all $t\in[T]$, the function $h_{t}:\mathbb{R}\times\mathbb{R}^{T+1}\times\mathbb{R}^{n^{\text{x}}}\rightarrow\mathbb{R}$ is a ${{\cal C}{\cal P}}(1)$ function.
@@
Assumption 8 comprises a collection of regularity conditions on the model attributes. The first two parts ensure that the functions $g_{t}$ and $h_{t}$ do not demonstrate fast explosive behavior, thereby ensuring the well-posedness of the large system asymptotic. The final part pertains to the initial observation of the network and corroborates that the function $g_{0}$ is non-degenerate. This ensures that initial observations provide meaningful information and contribute to the evolution of outcomes. {assumption} Fix $T\in\mathbb{N}$, $k\geq 2$, and a subpopulation $\mathcal{S}\subset[N]$, where the size $\left|{\mathcal{S}}\right|$ grows to infinity as $N\rightarrow\infty$. We assume the following statements hold:
@@
    Let $p_{y}^{\mathcal{S}}(N)$ denote the empirical distribution of the initial outcomes $Y^{i}_{0}{}$ with $i\in\mathcal{S}$, in a system with $N$ units. Then, $p_{y}^{\mathcal{S}}(N)$ converges weakly to a probability measure $p_{y}^{\mathcal{S}}$.
@@
    Let $p_{x}^{\mathcal{S}}(N)$ denote the empirical distribution of the covariate vectors ${\vec{X}}^{i}$ with $i\in\mathcal{S}$, in a system with $N$ units. Then, $p_{x}^{\mathcal{S}}(N)$ converges weakly to a probability measure $p_{x}^{\mathcal{S}}$.
@@
    Let $p_{w}^{\mathcal{S}}(N)$ denote the empirical distribution of the treatment vectors $\vec{W}^{i}$ with $i\in\mathcal{S}$, in a system with $N$ units. Then, $p_{w}^{\mathcal{S}}(N)$ converges weakly to a probability measure $p_{w}^{\mathcal{S}}$.
@@
    For all $i$ and any $t\in[T]_{0}$, let $p_{\mu^{i}}(N)$ and $p_{\mu_{t}^{i}}(N)$ be, respectively, the empirical distribution of the elements of the vectors $\vec{\mu}^{\;i\cdot}:=(\mu^{i1},\ldots,\mu^{iN})^{\top}$ and $\vec{\mu}^{\;i\cdot}_{t}:=(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})^{\top}$. Then, $p_{\mu^{i}}(N)$ converges weakly to $p_{\mu^{i}}$ and $p_{\mu_{t}^{i}}(N)$ converges weakly to $p_{\mu_{t}^{i}}$.